OpenStudy (anonymous):

Solve the system using elimination. {-8-6x=-7y 1/2=-x+3/4y Please help!

OpenStudy (mathstudent55):

First, rewrite both equations in the form ax + by = c. Can you do that?

OpenStudy (anonymous):

yeah i can do that, -8=6x-7y 1/2=-x+ 3/4 y

OpenStudy (anonymous):

so then i multiplied the second equation by 6 to elminate the "x" but I keep getting lost, am i even in the right ball park

OpenStudy (mathstudent55):

Great. Let's just switch sides on both equations. \( 6x - 7y = -8\) \(-x + \dfrac{3}{4}y = \dfrac{1}{2} \)

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

So then its -6x+4.5y=3 then I subtract from orginal equation to get -2.5y=-5, y=2 ???

OpenStudy (mathstudent55):

\( 6x - 7y =8\) \(-6x + 4.5x = 3\)

OpenStudy (mathstudent55):

Now add the equations together. 2.5x = 5 x = 2

OpenStudy (anonymous):

okay then u plug x into first equation?

OpenStudy (mathstudent55):

Yes.

OpenStudy (anonymous):

6x-7y=8 6(2)-7y=8 12-7y=8 -7y=-4 y=4/7 or .571 ??? So x=2 and y = 4/7 ???

OpenStudy (mathstudent55):

Wait. There's a mistake from before.

OpenStudy (anonymous):

yeah thought so, can't answer this question for the life of me!!!

OpenStudy (mathstudent55):

-8-6x=-7y ------> 6x - 7y = -8 (remember the -8, not 8) 1/2=-x+3/4y ------> -x + 0.75y = 0.5 Now we muiltiply tghe second equation by 6, and keep the first one the same: 6x - 7y = -8 -6x + 4.5y = 3 ----------------------(add) -2.5y = -5 y = 2

OpenStudy (mathstudent55):

Now we substitute x = 2 into the first equation: 6x - 7(2) = -8 6x - 14 = -8 6x = 6 x = 1

OpenStudy (anonymous):

so y=1 and x=2 ?

OpenStudy (anonymous):

or wait its y=2 and x=1 !

OpenStudy (anonymous):

Right mathstudent55? that y=2 and x=1

OpenStudy (mathstudent55):

Correct. x = 1 and y = 2

OpenStudy (anonymous):

Thank you so much, kept messing up this problem!

OpenStudy (mathstudent55):

You're welcome.